# Enhancing the Harrow-Hassidim-Lloyd (HHL) algorithm in systems with large condition numbers

### v2

@@ Acknowledgements.
We also present these results for $\alpha$ set to 60, 70, and 80 degrees (the main text only discussed the first case). Furthermore, while the main text considers 1 million shots, we also perform the same analyses for $10^{5}$ shots. All of these results are presented in Figures A2 and A3 for the cases considering $10^{5}$ and $10^{6}$ shots respectively. The accompanying data is provided in Tables A1, A2, and A3, for 60, 70, and 80 degrees examples respectively.
@@
Table A1: Table presenting the data for HHL and Psi-HHL algorithms (with $\alpha=60^{\circ}$) for $4\times 4$ matrices, and in particular, for three cases: $A$ diagonal and $\vec{b}$ in equal superposition (denoted in the table as ‘$A$ diag, $\vec{b}$ equal’), $A$ diagonal and $\vec{b}$ not in equal superposition (denoted in the table as ‘$A$ diag, $\vec{b}$ unequal’), and $A$ not diagonal and $\vec{b}$ in an unequal superposition (denoted in the table as ‘$A$ not diag, $\vec{b}$ unequal’). The string ‘nan’ refers to situations where we get the square root of a negative quantity for energy, which is unphysical, and hence is not a valid result.
@@
Table A2: Table providing the data for HHL and Psi-HHL algorithms (with $\alpha=70^{\circ}$) for $4\times 4$ matrices. The notations are the same as those used in Table A1.
@@
Table A3: Table providing the data for HHL and Psi-HHL algorithms (with $\alpha=80^{\circ}$) for $4\times 4$ matrices. The notations are the same as those used in Tables A1 and A2.
@@
Table A4: Summary of the fitted probabilities of HHL and Psi-HHL for the 4x4 case $A$ not diagonal and $\vec{b}$ in an unequal superposition of states, $\alpha=60^{\circ},70^{\circ},80^{\circ}$ and the number of shots is fixed to $10^{6}$.
@@
Figure A4: Sub-figures presenting the average probabilities of HHL and Psi-HHL with varying condition number $\mathcal{\kappa}$ of $A$, for $8\times 8$ matrices using $10^{5}$ shots ((a) through (c)) and using $10^{6}$ ((d) through (f))
@@
Figure A5: Sub-figures presenting the average probabilities of HHL and Psi-HHL with varying condition number $\mathcal{\kappa}$ of $A$, for $16\times 16$ matrices and using $10^{6}$ shots throughout.
@@
Figure A6: Data for $32\times 32$ matrices. Note that we have again reproduced here the sub-figure (a) with 1 million shots (which are already present in the main text).
@@
Figure A7: Data for $64\times 64$ matrices. Note that we have again reproduced here the sub-figure (a) with 1 million shots (which are already present in the main text).
Figure A8: Figure showing average (over 10 repetitions) values of percentage fraction difference (PFD) of HHL and Psi-HHL, as the condition number grows from $2^{2}$ to $2^{17}$, for fixed $A$ matrix sizes going from $4\times 4$ to $64\times 64$. Only the case where $\alpha=60^{\circ}$, $A$ not diagonal, $\vec{b}$ in an unequal superposition, and number of shots = $10^{6}$ is considered. The blue cells indicate cases where HHL gives outputs, and red denotes cases where HHL fails. The white cells denote those cases for which we did not carry a calculation, due to computational cost.
Table A5: Table providing the data for HHL and Psi-HHL algorithms (with $\alpha=60^{\circ}$, $70^{\circ}$, $80^{\circ}$) for $8\times 8$ matrices.
@@
Table A6: Table providing the data for HHL and Psi-HHL algorithms (with $\alpha=60^{\circ}$, $70^{\circ}$, $80^{\circ}$) for $16\times 16$ matrices.
@@
Table A7: Table providing the data for HHL and Psi-HHL algorithms (with $\alpha=60^{\circ}$, $70^{\circ}$, $80^{\circ}$) for $32\times 32$ matrices.
@@
Table A8: Table providing the data for HHL and Psi-HHL algorithms (with $\alpha=60^{\circ}$, $70^{\circ}$, $80^{\circ}$) for $64\times 64$ matrices.
@@
We begin by explaining the procedure for generating the $A$ matrix of size $n$. We use the configuration state function-based Graphical Unitary Group Approach-Configuration Interaction Singles and Doubles (GUGA-CISD) method available in the GAMESS program [27] to construct the Hamiltonian for a given molecule in a specified single particle basis. GUGA-CI ensures that the Hamiltonian is constructed in an orthogonal basis; hence, it is Hermitian. Also, because it is a spin-adapted CI method, it avoids spin contamination of the wave function, thereby giving accurate results. We remove the first row and the first column of the GUGA-CISD Hamiltonian matrix and subtract the diagonal entries of the resulting matrix with the energy obtained from Møller-Plesset theory to second order in perturbation (also called MP2 in literature). The latter step is carried out for the purpose of normal ordering, which in turn ensures that HHL would output the correlation part of the ground state energy directly instead of the total ground state energy (which can be written as a sum of the HF energy and the correlation energy, with the latter being much lesser than the former). This is desirable when one executes HHL-LCC calculations on NISQ era quantum hardware, as the error in the total ground state energy can be much greater than the correlation energy itself. We also note that unlike the traditional approach where the HF energy is subtracted from the diagonal entries, we take away the MP2 energy, in order to make sure that $A$ is positive definite. In a sense, subtracting MP2 energy could be thought of as a level shift over and beyond subtracting the HF energy.
@@
We now focus our attention on the elements of the $A$ matrix. The size of an $A$ matrix, $m$, generated from a quantum chemical calculation, which we can denote as $A_{mol}$, is not necessarily of the dimension $n\times n$. In such cases, we embed the $A_{mol}$ matrix of dimension ${m\times m}$ as a diagonal block in a larger matrix of dimension $n\times n$, with the other diagonal block being an identity matrix of size $(n-m)\times(n-m)$, denoted as $\mathcal{I}$. In such a case, we solve a system of linear equations given by
$A=\begin{bmatrix}A_{mol}&0\\ 0&\mathcal{I}\end{bmatrix}\begin{bmatrix}\vec{x}\\ 0\end{bmatrix}=\begin{bmatrix}\vec{b}\\ 0\end{bmatrix}.$ (A25)
@@
This approach ensures that the eigenvalues of $A_{mol}$ remain unaffected, as long as its maximum eigenvalue is $\approx$ 1. We choose our matrices carefully so that this condition is satisfied. In the event that the maximum diagonal entry of $A$ is less than 1, it is recommended to pad with the largest (or the smallest) diagonal entry of $A$ so that the number of shots expended is reduced. As an example, we picked a $3\times 3$ sparse matrix such that the largest diagonal entry is 0.70: $\begin{pmatrix}0.25&0.10&0.00\\ 0.10&0.45&0.20\\ 0.00&0.20&0.70\end{pmatrix}$, and considered two cases: padding with 1 and with 0.70. We found that kappa is larger in the latter (5.16) than in the former (4.21).
@@
Figure A9: Sub-figure (a) shows the increase in system size with number of virtuals, $n_{v}$, for KH, RbH and CsH molecules in the LCCSD approximation, while the sub-figures (b), (c) and (d) show that the $A$ matrix of respectively KH, RbH and CsH increases as $n_{o}^{2}n_{v}^{2}$ for both cases, as one would expect.
@@
We fix the number of core orbitals to 6, 8 and 10. We now describe the method adopted in choosing virtual molecular orbitals (MOs). Instead of taking the approach of ordering the virtual MOs by energy in ascending order and then imposing an energy cut-off to define our active space, we pick those MOs that are dominated by important atomic orbitals (AOs). We select the important AOs, and for the purposes of this work, we pick those AOs in the order of the principal quantum numbers. For example, in the case of RbH, we can decide the important AOs from the set that arise from our choice of basis, namely the Sapporo double zeta basis: $5s$ through $7p$. For the purposes of our analysis on condition number versus system size, we pick all the AOs that the basis allows in order to obtain more data points for fitting (details given in Table A9). We now move to some specifics for RbH: for the smallest system size considered (1 virtual MO), we pick the first virtual from our sorted MO list, which is the MO dominated by the Rb $5s-5p_{z}$ hybrid, and thus our active space that is built out of 6 occupied MOs and 1 virtual one has a system size, $m$, of $209\times 209$, and subsequently upon padding, an $A$ matrix size, $n$, of $256\times 256$. For the next system size, we add the next MO in the sorted list, which is dominated by Rb $5s$, leading to a $279\times 279$ matrix, which, after padding, becomes a matrix of size $512\times 512$. We continue to construct $A$ matrices in this fashion until we reach $m=2296$, for which $\mathcal{\kappa}$ is about 389.76.
@@
Table A9: Table providing the data for condition number ($\mathcal{\kappa}$) of the $m\times m$ $A$ matrix versus system size ($m$), for KH, RbH and CsH molecules, where we vary the virtual orbitals to increase $m$. In the table, $n_{o}$ is the number of active spin orbitals, and $n_{v}$ is the number of unoccupied spin orbitals. The column labelled ‘Reordered virtual MOs’ represents the atomic orbitals reordered according to the ones that contribute dominantly to the virtual molecular orbital. The column labelled ‘MO Energies’ gives the molecular virtual orbital energies in units of milli-Hartree.
@@
Table A11: Table providing the minimum and maximum eigenvalues of the matrix $\mathrm{A}_{mol}$ of size $m$, denoted in the table as $\lambda_{\mathrm{min}}^{\mathrm{A_{mol}}}$ and $\lambda_{\mathrm{max}}^{\mathrm{A_{mol}}}$ respectively, for KH, RbH and CsH systems. $n_{o}$ and $n_{v}$ denote the number of occupied and virtual orbitals respectively. The energies are in units of milli-Hartree.
@@
m $n_{v}$ $\lambda_{\mathrm{min}}^{\mathrm{A_{mol}}}$ $\lambda_{\mathrm{max}}^{\mathrm{A_{mol}}}$ m $n_{v}$ $\lambda_{\mathrm{min}}^{\mathrm{A_{mol}}}$ $\lambda_{\mathrm{max}}^{\mathrm{A_{mol}}}$ m $n_{v}$ $\lambda_{\mathrm{min}}^{\mathrm{A_{mol}}}$ $\lambda_{\mathrm{max}}^{\mathrm{A_{mol}}}$
@@
Table A12: Average PFDs and Range of PFDs of HHL and Psi-HHL for Li2 and KH molecules, both in Sapporo-DKH3-DZP-2012 basis set. The PFDs are calculated over a range of 30 repetitions.
@@
Table A13: Table providing the details on the condition numbers ($\kappa$), the average total energy, the energy difference ($\Delta E$) between the total HHL/Psi-HHL energy and the LCCSD total energy and the PFDs over 30 repetitions with respect to the bond length of KH molecule. All the energies are in units of milli-Hartree.

### v3

@@ Acknowledgements.
We also present these results for $\alpha$ set to 60, 70, and 80 degrees (the main text only discusses the first case). Furthermore, while the main text considers 1 million shots, we also perform the same analyses for $10^{5}$ shots. All of these results are presented in Figures A4 and A5 for the cases considering $10^{5}$ and $10^{6}$ shots respectively. The accompanying data is provided in Tables A4, A5, and A6, for 60, 70, and 80 degrees examples respectively.
@@
Table A4: Table presenting the data for HHL and Psi-HHL algorithms (with $\alpha=60^{\circ}$) for $4\times 4$ matrices, and in particular, for three cases: $A$ diagonal and $\vec{b}$ in equal superposition (denoted in the table as ‘$A$ diag, $\vec{b}$ equal’), $A$ diagonal and $\vec{b}$ not in equal superposition (denoted in the table as ‘$A$ diag, $\vec{b}$ unequal’), and $A$ not diagonal and $\vec{b}$ in an unequal superposition (denoted in the table as ‘$A$ not diag, $\vec{b}$ unequal’). The string ‘nan’ refers to situations where we get the square root of a negative quantity for energy, which is unphysical, and hence is not a valid result. Each data point is generated by taking the average over 50 repetitions.
@@
Table A5: Table providing the data for HHL and Psi-HHL algorithms (with $\alpha=70^{\circ}$) for $4\times 4$ matrices. The notations are the same as those used in Table A4. Each data point is generated by taking the average over 50 repetitions.
@@
Table A6: Table providing the data for HHL and Psi-HHL algorithms (with $\alpha=80^{\circ}$) for $4\times 4$ matrices. The notations are the same as those used in Tables A4 and A5. Each data point is generated by taking the average over 50 repetitions.
@@
Table A7: Summary of the fitted probabilities obtained from HHL and Psi-HHL for the 4x4 case of $A$ not diagonal and $\vec{b}$ in an unequal superposition, for $\alpha=60^{\circ},70^{\circ},80^{\circ}$. The number of shots is fixed to $10^{6}$.
@@
Figure A6: Sub-figures presenting the average probabilities, PFDs and range of PFDs of HHL and Psi-HHL with varying condition number $\mathcal{\kappa}$ of $A$, for $8\times 8$ matrices using $10^{5}$ shots ((a) through (c)) and using $10^{6}$ ((d) through (f)). In all of the examples, $A$ is chosen to be a non-diagonal matrix and $\vec{b}$ is set in an unequal superposition. Each data point is an average over 10 repetitions.
@@
Figure A7: Sub-figures presenting the average probabilities, PFDs and range of PFDs of HHL and Psi-HHL with varying condition number $\mathcal{\kappa}$ of $A$, for $16\times 16$ matrices and using $10^{6}$ shots throughout. In all of the examples, $A$ is chosen to be a non-diagonal matrix and $\vec{b}$ is set in an unequal superposition. Each data point is an average over 10 repetitions.
@@
Figure A8: Data for $32\times 32$ matrices. Note that we have again reproduced here the sub-figure (a) with 1 million shots (which are already present in the main text). In all of the examples, $A$ is chosen to be a non-diagonal matrix and $\vec{b}$ is set in an unequal superposition. Each data point is an average over 10 repetitions.
@@
Figure A9: Data for $64\times 64$ matrices. Note that we have again reproduced here the sub-figure (a) with 1 million shots (which are already present in the main text). In all of the examples, $A$ is chosen to be a non-diagonal matrix and $\vec{b}$ is set in an unequal superposition. Each data point is an average over 10 repetitions.
Figure A10: Figure showing average (over 10 repetitions) values of percentage fraction difference (PFD) obtained from HHL and Psi-HHL, as the condition number, $\mathcal{\kappa}$, grows from $2^{2}$ to $2^{20}$, for fixed $A$ matrix sizes going from $4\times 4$ to $64\times 64$. Only the case where $\alpha=60^{\circ}$, $A$ not diagonal, $\vec{b}$ in an unequal superposition, and number of shots = $10^{6}$ is considered. The blue cells indicate cases where HHL gives outputs, and red denotes cases where HHL fails. The white cells denote those cases for which we did not carry out a calculation, due to computational cost.
Table A8: Table providing the data for HHL and Psi-HHL algorithms (with $\alpha=60^{\circ}$, $70^{\circ}$, $80^{\circ}$) for $8\times 8$ matrices. $A$ is a non-diagonal matrix and $\vec{b}$ is an unequal superposition. Each data point is an average over 10 repetitions.
@@
Table A9: Table providing the data for HHL and Psi-HHL algorithms (with $\alpha=60^{\circ}$, $70^{\circ}$, $80^{\circ}$) for $16\times 16$ matrices. $A$ is a non-diagonal matrix and $\vec{b}$ is an unequal superposition. Each data point is an average over 10 repetitions.
@@
Table A10: Table providing the data for HHL and Psi-HHL algorithms (with $\alpha=60^{\circ}$, $70^{\circ}$, $80^{\circ}$) for $32\times 32$ matrices. $A$ is a non-diagonal matrix and $\vec{b}$ is an unequal superposition. Each data point is an average over 10 repetitions.
@@
Table A11: Table providing the data for HHL and Psi-HHL algorithms (with $\alpha=60^{\circ}$, $70^{\circ}$, $80^{\circ}$) for $64\times 64$ matrices. $A$ is a non-diagonal matrix and $\vec{b}$ is an unequal superposition. Each data point is an average over 10 repetitions.
@@
We begin by explaining the procedure for generating an $A$ matrix of size $n\times n$. We use the configuration state function-based Graphical Unitary Group Approach-Configuration Interaction Singles and Doubles (GUGA-CISD) method available in the GAMESS program [27] to construct the Hamiltonian for a given molecule in a specified single particle basis. GUGA-CI ensures that the Hamiltonian is constructed in an orthogonal basis; hence, it is Hermitian. Also, because it is a spin-adapted CI method, it avoids spin contamination of the wave function, thereby giving accurate results. In order to generate the $A$ matrix, we remove the first row and the first column of the GUGA-CISD Hamiltonian matrix and subtract the diagonal entries of the resulting matrix with the energy obtained from Møller-Plesset theory to second order in perturbation (also called MP2 in literature). The latter step is carried out for the purpose of normal ordering, which in turn ensures that HHL would output the correlation part of the ground state energy directly instead of the total ground state energy (which can be written as a sum of the HF energy and the correlation energy, with the latter being much smaller than the former). This is desirable when one executes HHL-LCC calculations on NISQ/late-NISQ era quantum hardware, as the error in the total ground state energy can be much greater than the correlation energy itself. We also note that unlike the traditional approach where the HF energy is subtracted from the diagonal entries, we take away the MP2 energy, in order to make sure that $A$ is positive definite. In a sense, subtracting MP2 energy could be thought of as a level shift over and beyond subtracting the HF energy.
@@
We now focus our attention on the elements of the $A$ matrix. The size of an $A$ matrix, $m\times m$, generated from a quantum chemical calculation, which we can denote as $A_{mol}$, is not necessarily of the dimension $n\times n$. In such cases, we embed the $A_{mol}$ matrix of dimension ${m\times m}$ as a diagonal block in a larger matrix of dimension $n\times n$, with the other diagonal block being an identity matrix of size $(n-m)\times(n-m)$, denoted as $\mathcal{I}$. In such a case, we solve a system of linear equations given by
$A=\begin{pmatrix}A_{mol}&0\\ 0&\mathcal{I}\end{pmatrix}\begin{pmatrix}\vec{x}\\ \vec{0}\end{pmatrix}=\begin{pmatrix}\vec{b}\\ \vec{0}\end{pmatrix}.$ (A25)
@@
This approach ensures that the eigenvalues of $A_{mol}$ remain unaffected, as long as its maximum eigenvalue of the scaled matrix is $\approx$ 1. We choose our matrices carefully so that this condition is satisfied. In the event that the maximum diagonal entry of $A$ is less than 1, it is recommended to pad with the largest (or the smallest) diagonal entry of $A$ so that the number of shots expended is reduced. As an example, we picked a $3\times 3$ sparse matrix such that the largest diagonal entry is 0.70: $\begin{pmatrix}0.25&0.10&0.00\\ 0.10&0.45&0.20\\ 0.00&0.20&0.70\end{pmatrix}$, and considered two cases: padding with 1 and with 0.70. We found that the condition number, $\mathcal{\kappa}$, is larger in the latter (5.16) than in the former (4.21).
@@
Figure A11: Sub-figure (a) shows the increase in system size ($m$) with the number of virtuals, $n_{v}$, for KH, RbH and CsH molecules in the LCCSD approximation, while the sub-figures (b), (c) and (d) show that $m$ increases as $n_{o}^{2}n_{v}^{2}$ for the three molecules, as one would expect.
@@
We fix the number of occupied orbitals to 6, 8 and 10. We now describe the method adopted in choosing virtual molecular orbitals (MOs). Instead of taking the approach of ordering the virtual MOs by energy in ascending order and then imposing an energy cut-off to define our active space, we pick those MOs that are dominated by important atomic orbitals (AOs). We select the important AOs, and for the purposes of this work, we pick those AOs in the order of the principal quantum numbers. For example, in the case of RbH, we can decide the important AOs from the set that arise from our choice of basis, namely the Sapporo double zeta basis: $5s$ through $7p$. For the purposes of our analysis on condition number versus system size, we pick all the AOs that the basis allows in order to obtain more data points for fitting (details given in Table A12). We now move to some specifics for RbH: for the smallest system size considered (1 virtual MO), we pick the first virtual from our sorted MO list, which is the MO dominated by the Rb $5s-5p_{z}$ hybrid, and thus our active space that is built out of 6 occupied MOs and 1 virtual one has a system size, $m\times m$, of $209\times 209$, and subsequently upon padding, an $A$ matrix size, $n\times n$, of $256\times 256$. For the next system size, we add the next MO in the sorted list, which is dominated by Rb $5s$, leading to a $279\times 279$ matrix, which, after padding, becomes a matrix of size $512\times 512$. We continue to construct $A$ matrices in this fashion until we reach $m=2296$, for which $\mathcal{\kappa}$ is about 211.94.
@@
Table A12: Table providing the data for condition number ($\mathcal{\kappa}$) of the $m\times m$ $A_{mol}$ matrix versus system size ($m$), for KH, RbH and CsH molecules, where we vary the virtual orbitals to increase $m$. In the table, $n_{o}$ is the number of active spin orbitals, and $n_{v}$ is the number of unoccupied spin orbitals. The column labelled ‘Reordered virtual MOs’ represents the atomic orbitals reordered according to the ones that contribute dominantly to the virtual molecular orbital. The column labelled ‘MO energies’ gives the virtual molecular orbital energies in units of milli-Hartree.
@@
Table A14: Table providing the minimum and maximum eigenvalues of the matrix $A_{mol}$ of size $m\times m$, denoted in the table as $\lambda_{\mathrm{min}}^{A_{mol}}$ and $\lambda_{\mathrm{max}}^{A_{mol}}$ respectively, for KH, RbH and CsH systems. $n_{o}$ and $n_{v}$ denote the number of occupied and virtual orbitals, respectively. $\lambda_{\mathrm{min}}^{A_{mol}}$ and $\lambda_{\mathrm{max}}^{A_{mol}}$ are in units of milli-Hartree.
@@
m $n_{v}$ $\lambda_{\mathrm{min}}^{A_{mol}}$ $\lambda_{\mathrm{max}}^{A_{mol}}$ m $n_{v}$ $\lambda_{\mathrm{min}}^{A_{mol}}$ $\lambda_{\mathrm{max}}^{A_{mol}}$ m $n_{v}$ $\lambda_{\mathrm{min}}^{A_{mol}}$ $\lambda_{\mathrm{max}}^{A_{mol}}$
@@
Table A15: Average PFDs and range of PFDs of HHL and Psi-HHL for Li2 and KH molecules, both in Sapporo-DKH3-DZP-2012 basis set. The average PFDs are calculated over a range of 30 repetitions.
@@
Table A16: Table providing the details on the condition numbers ($\kappa$), the average total energy, the energy difference ($\Delta E$) between the total HHL (Psi-HHL) energy and classical value of the LCCSD total energy and the average PFDs over 30 repetitions with respect to the bond length of KH molecule. All the energies are in units of milli-Hartree. ‘Eq’ denotes the equilibrium geometry.